Bobbiwasabi007:

A line containing the points $(9, 1)$ and $(5, 5)$ intersects the $x$ -axis at what point?

Mercury:

we want to write the equation between these two points first apply the slope formula first m = (y2-y1)/(x2-x1) then once you have m, take either of your points to find b, using y = mx + b once you have your m and b values, create your y = mx + b equation. now, since we want the value at which the graph intersects the x-axis, let y = 0 and solve for x